OpenStudy (honeymonster):

can you help me with this question please..equation posted in replies..find the value of x

OpenStudy (honeymonster):

\[3-4 (x-1)=0\]

OpenStudy (kropot72):

Firstly you need to multiply out to remove the brackets: -4(x - 1) = -4x + 4 Do you follow this?

OpenStudy (cathyangs):

There are several ways to solve this type of equation. First, follow the order of operations to simplify the equation as much as you can.

OpenStudy (kropot72):

Then you need to state the problem as : 3 - 4x + 4 = 0

OpenStudy (cathyangs):

Oops XD didn't see that you've got this, kropot. Sorry.